OpenStudy (anonymous):

A state has $5,000,000 worth of bonds that are due in 20 years. A sinking fund is established to pay off the debt. If the state can earn 10% annually on it's money, what is the annual sinking fund deposit needed? Prepare a table showing the growth of the sinking fund every 5 years. (Please Help)

OpenStudy (jack1):

not that good with interest rate formulas, but i think: debt = deposit × (1+interest rate)^(number of periods) so 5000000 = deposit × (1+0.10)^(20) \[5,000,000 = deposit \times (1.10)^{20}\] give that a shot dude, and tell me how u go?

OpenStudy (anonymous):

ok..thanks alot...jst tried

OpenStudy (anonymous):

it correct Jack 1...the answer is $87,298.13 is the value of deposit so in table format for 5 years.

OpenStudy (jack1):

sweet, glad it worked @PeceliVura !

OpenStudy (jack1):

wait, ... what? i got deposit = 743000 ish

OpenStudy (jack1):

how did u get ur answer?

OpenStudy (anonymous):

ok let me show my work...

OpenStudy (anonymous):

5000000=A[(1+0.10)^20-1]/0.10 5000000=A(57.27499949) 5000000/57.27499949=A $87,298.13=A

OpenStudy (anonymous):

does it right...

OpenStudy (jack1):

ummm... i'm not sure where ur formula came from, sorry i don;t know if my way is right or not, but i did it: FV=PV×(1+r)^n FV = future Value PV = present value r = interest rate n = number of periods 5000000=deposit×(1+0.10)^20 5000000 = deposit x (1.1)^20 5000000 = doposit x 6.7275 5000000/6.7275 = deposit deposit = $743,218

OpenStudy (anonymous):

Jack that is the Ordinary annuity...FV=A[(1+i)^n-1/i] what do you think which one to used...pls

OpenStudy (jack1):

ahhh, sorry i see it now, annual deposit!

OpenStudy (anonymous):

thank you so my sweet...god bless you and you famly

OpenStudy (jack1):

sorry, mabad

OpenStudy (jack1):

\[FV=\frac{A(1+i)^n-1}i \rightarrow \rightarrow or \rightarrow \rightarrow FV= A[\frac {(1+i)^{n-1}}i]\] is ur equation the frist one or the seconed one?

OpenStudy (anonymous):

first one..

OpenStudy (anonymous):

can I asked you another question please....sweet.

OpenStudy (anonymous):

I have an assigment which was due today Friday but I will submit on Monday...coz I need help more

OpenStudy (jack1):

$87,298.10, ur spot on then above, sorry for my basic answer before

OpenStudy (jack1):

of course ask away, i'll help if i can ;)

OpenStudy (anonymous):

no I take your answer...I belive it is right...

OpenStudy (anonymous):

wait...sorry

OpenStudy (anonymous):

u have email address...to up load question.

OpenStudy (jack1):

sorry peli, i gave it out to strangers once and hot hacked, had to get a new one u can post stuff by hitting the attach file button below the here, its blue

OpenStudy (anonymous):

ok sorry...thanks

OpenStudy (anonymous):

Attached are detail of my assignment...

OpenStudy (anonymous):

I am finished with Q1,Q8,Q5 and Q2....please can you help me on this Jack

OpenStudy (jack1):

sure so Q3 a is easy enough just add the cells downwards

OpenStudy (jack1):

so Plan 1 10+11+11+12+12+13= nd Plan 2 10.3 + 10.6+10.9+11.2+11.5+11.8=

OpenStudy (anonymous):

ok

OpenStudy (jack1):

sp plan 1 = 9000 better than just base salary of 10000 nd plan 2 = 6300 better than just base salary of 10000

OpenStudy (jack1):

so plan 1 is better and by 2700 (for a 3 yr period)

OpenStudy (jack1):

next for 5 yrs so would be: Plan 1: 10+11+11+12+12+13 + 13 + 14 + 14 + 15= Plan 2: 10.3 + 10.6+10.9+11.2+11.5+11.8 + 12.1 + 12.4+ 12.7 +13.0=

OpenStudy (anonymous):

so how u get the 9000 pls

OpenStudy (jack1):

10+11+11+12+12+13= 69000 -60000 if no raise = benefit of $9000 over a 3 yr period

OpenStudy (anonymous):

ok...thanks..

OpenStudy (anonymous):

where you get 60000 if no raise

OpenStudy (jack1):

so plan 1, 5 yrs = 125000 -100000 if no raise = benefit of 25000 over 5 yrs so plan 2 = 116500 -100000 if no raise = benefit of 16500 over 5 years

OpenStudy (jack1):

base salary is $10000 per 6 months, and we're comparing the 2 optional raises over a 3 yr period, at just the base salary, no raises, u'd earn $60000 over a 3 yr period, with option 1 (+1000/yr), u'd earn $69000 over a 3 yr period, with option 2 (+300/6months), u'd earn $66300

OpenStudy (anonymous):

ok...you are so brain box...so smart..thanks alot

OpenStudy (jack1):

all good, n e time

OpenStudy (anonymous):

hahaha...ok

OpenStudy (jack1):

so again, over 5 yrs, plan 1 rocks better than plan 2

OpenStudy (jack1):

25000 vs 16500 (5 yrs)

OpenStudy (anonymous):

that is quetion 3. I have load it into my flash than re-write it and submit...it a big help you have done..thanks

OpenStudy (jack1):

all good, r u ok with q 6?

OpenStudy (anonymous):

but practiced make perfect

OpenStudy (jack1):

no worries, good luck

OpenStudy (anonymous):

no please...

OpenStudy (anonymous):

My answer Q6.A is $2,028,758.14 Formular used FV=A[(1+i)6n -1]/i

OpenStudy (anonymous):

B. $1,022,802.61 same formula

OpenStudy (anonymous):

does it right sweet...

OpenStudy (jack1):

i gots roof equals 180611 in 20 yrs time...?

OpenStudy (jack1):

i used the formula i did whay up above tho...?

OpenStudy (jack1):

g2g diner, back later ? soz

OpenStudy (jack1):

hey, im back k i still think that the cost of the roof (one off payment) is $180,611 using FV=PV×(1+r)^n so that's what i got for 6a anyway 6b is kinda dependant on what u get for 6a, but

OpenStudy (jack1):

cool dude, good to c u hey ;)

OpenStudy (jack1):

so 6b i used your formula from whay up above and calculated it out at : 180611 = x ((((1.015)^40)-1)/.015) so x = $3,328.14 as a semiannual payment

OpenStudy (jack1):

FV=A[ ( {(1+i)^n} - 1 ) /i ] this was the equation i used to work it out using i = 0.015 n = 40 and FV = 180611

OpenStudy (jack1):

again, i don't know if its correct dude, i'm rusty when it comes to compound interest formulas hay

OpenStudy (jack1):

dammmit, sorry the annual interest rate on treasuries is 5%, not 3 used the wrong one, standby im recalculating

OpenStudy (jack1):

FV=A[ ( {(1+i)^n} - 1 ) /i ] this was the equation i used to work it out using i = 0.025 n = 40 and FV = $180, 611 so 180611 = x ((((1.025)^40)-1)/0.025) 180611 = x (67.4026) so x = 180611/67.4026 x = 2679.59 = A

OpenStudy (jack1):

yep, a semi-annual payment of $2,679.59 will reach the goal in 20 yrs

OpenStudy (jack1):

so 6a = 180611 6b = 2,679.59

OpenStudy (jack1):

all good dude, good luck with the rest think u've got Q1, Q2, Q3, Q5, Q6, Q8, , just 4, 7 and 9 to go

OpenStudy (anonymous):

I have Q4, but undifind...my formula is r=(1+i/n)-1 just check